OpenStudy (anonymous):

solve -2/5x-9< 9/10 show your steps thanks for the help

OpenStudy (anonymous):

anybody? any help is appreciated

OpenStudy (anonymous):

\[-\frac{ 2 }{ 5x-9 }<\frac{ 9 }{ 10 }\]

OpenStudy (anonymous):

is this your question?

OpenStudy (anonymous):

no is -2 over 5 times x -9< 9 over 10

OpenStudy (anonymous):

its*

OpenStudy (anonymous):

\[\frac{ -2 }{ 5(x-9) }<\frac{ 9 }{ 10 }\]

OpenStudy (anonymous):

that?

OpenStudy (anonymous):

Use the equation button to make it if I got it wrong again.

OpenStudy (anonymous):

here let me draw it

OpenStudy (anonymous):

\[-2/5 \bullet x -9<9/10\]

OpenStudy (anonymous):

someobdy help plz?

OpenStudy (anonymous):

ooh

OpenStudy (anonymous):

\[\frac{ -2 }{ 5 }\times x-9 <\frac{ 9 }{ 10 }\]

OpenStudy (anonymous):

x > 27/4

OpenStudy (anonymous):

cool , thanks , but could you show your steps?

OpenStudy (anonymous):

\[-2/5(x-9) = -2/5+18/5 < 9/10\]

OpenStudy (anonymous):

first step

OpenStudy (anonymous):

it should be -2/5x sorry

OpenStudy (anonymous):

\[-2/5x < 9/10 -18/5 => -2/5x <-27/10\]

OpenStudy (anonymous):

\[x > -27/10 \times -5/2 = > x = 27/4\]

OpenStudy (anonymous):

thanks I got it and your a lifesaver

OpenStudy (anonymous):

sign changes because we divided by negative, and you're welcome